AH-453 kJ/mol-rxn P4010(S) + 6H₂O(1)-4H3PO4(aq) How many grams of P4010(S) would have to react to produce 167 kJ of energy?

1
Expert's answer
2022-05-31T03:25:03-0400

Mr (P4O10) = 284 g/mol


P4O10 + 6 H2O = 4 H3PO4 + 453 kJ

According to the above reaction, 453 kJ of energy is produced when 284 g of P4O10 reacts. We find how many grams of P4O10 need to react to produce 167 kJ of energy:

284 g P4O10 —> 453 kJ

X —> 167 kJ

X = 284 * 167 / 453 = 104,7 g

ANSWER: 104,7 g